OpenStudy (anonymous):

(1/1)+(5/a-1)=(7/6) solve and show work pls also how do you find the LCD?

OpenStudy (dumbcow):

combine fractions on left hand side to add these fractions denominators must be the same 1 and (a-1) have no common factors a-1 will be the LCD because 1*(a-1) = a-1 a-1/a-1 + 5/a-1 = 7/6 (a-1 +5)/a-1 = 7/6 a+4/a-1 = 7/6 we have 2 fractions equal to each other cross multiply 6*(a+4) = 7*(a-1) 6a + 24 = 7a -7 -a = -31 a = 31